# VTG-LLM: Integrating Timestamp Knowledge into Video LLMs for Enhanced Video Temporal Grounding

### v2

@@ Abstract
Video Temporal Grounding (VTG) focuses on accurately identifying event timestamps within a particular video based on a linguistic query, playing a vital role in downstream tasks such as video browsing and editing. While Video Large Language Models (video LLMs) have made significant progress in understanding video content, they often face challenges in accurately pinpointing timestamps within videos, which limits their performance on VTG tasks. Therefore, to improve video LLMs’ ability to effectively locate timestamps, we argue that two critical aspects need to be enhanced. First, it is essential to have high-quality instructional tuning datasets that encompass mainstream VTG tasks. Second, directly incorporating timestamp knowledge into video LLMs is crucial, as it enables models to efficiently comprehend timestamp information. To address these needs, we first introduce VTG-IT-120K, a high-quality and comprehensive instruction tuning dataset that covers VTG tasks such as moment retrieval, dense video captioning, video summarization, and video highlight detection. Furthermore, we propose a specially designed video LLM model for VTG tasks, VTG-LLM, which (1) effectively integrates timestamp knowledge into visual tokens; (2) incorporates absolute-time tokens that specifically handle timestamp knowledge, thereby avoiding concept shifts; and (3) introduces a lightweight, high-performance slot-based token compression method to facilitate the sampling of more video frames. Comprehensive experiments showcase the superior performance of VTG-LLM in comparison to other video LLM methods across various VTG tasks. Our code and datasets are available at https://github.com/gyxxyg/VTG-LLM.
1 Introduction
Large language models (LLMs) (Kaplan et al., 2020; Achiam et al., 2023; Touvron et al., 2023) have exhibited considerable potential in capturing knowledge and tackling real-world challenges using a zero-shot approach. Recently, research has explored integrating knowledge from other modalities, such as vision (Liu et al., 2024; Li et al., 2023a) and audio (Ghosal et al., 2023), to enhance the capabilities of LLMs. Within the visual modality, video large language models (video LLMs) have emerged as a significant research area (Lin et al., 2023a; Zhang et al., 2023; Li et al., 2023b). These models aim to generate textual outputs based on given videos and text prompts, achieving remarkable success in multiple video-related tasks like video captioning (Xu et al., 2024) and question-answering (Li et al., 2023c).
@@ 1 Introduction
Video Temporal Grounding (VTG) is a crucial component of video understanding. It requires models to accurately pinpoint event timestamps within a video based on the given query. This task is essential for subsequent operations such as video browsing and editing. The sub-tasks within VTG can be categorized based on output formats, as described by (Lin et al., 2023b). For instance, moment retrieval and dense video captioning (Fabian Caba Heilbron & Niebles, 2015; Zhou et al., 2018) tasks demand models to generate one or multiple timestamp intervals accompanied by captions; video highlight detection (Lei et al., 2021) task requires models to produce a salient score curve; and video summarization (Song et al., 2015) task calls for models to output a series of timestamps along with their corresponding video frames.
Despite the remarkable achievements of video LLMs in video understanding tasks, current models struggle to accurately pinpoint timestamps, leading to subpar performance on VTG tasks. Therefore, it is essential to enhance video LLMs by equipping them with the ability to effectively comprehend timestamps. To accomplish this, we contend that two key aspects need to be addressed. On one hand, high-quality instructional tuning datasets specifically tailored for VTG tasks are necessary. Existing datasets are either of low quality (Zellers et al., 2021) or exhibit significant task imbalance (Ren et al., 2023; Wang et al., 2024b). On the other hand, it is crucial to develop new models that efficiently and explicitly incorporate timestamp knowledge into video LLMs.
In this paper, we tackle the aforementioned issues by first introducing a high-quality instruction tuning dataset, VTG-IT-120K, which consists of 47.2K videos and 120K annotations. This dataset encompasses most mainstream VTG tasks, including moment retrieval (63.2K), dense video captioning (37.2K), video summarization (15.2K), and video highlight detection (3.9K). The VTG-IT-120K dataset is more task-balanced compared to existing datasets (Ren et al., 2023; Wang et al., 2024b), which either contain annotations exclusively for the moment retrieval task or provide a mere 75 samples for the video summarization task. Moreover, the VTG-IT-120K includes 51.9K self-collected data annotations of superior quality. Furthermore, we present a novel video LLM model, VTG-LLM, which comprises three components for efficiently integrating timestamp knowledge into video LLMs: (1) To ensure that the inputs of LLMs contain sufficient timestamp information, we introduce a sequence-time embedding method that integrates timestamp data into visual tokens; (2) Concept shifts can occur when different input data lead to the same output targets, which can obscure the decision boundary (Moreno-Torres et al., 2012). For instance, the number 20 could represent seconds or number of people. To mitigate this, we incorporate absolute-time tokens that specifically handle timestamp knowledge, without introducing quantization errors. This ensures that the LLM tokenizer and LLM prediction head do not confuse timestamp-related information with other digit-related information, such as counting and mathematics; (3) Given the limited context length of LLMs, the number of visual tokens must be reduced to allow for more video frame sampling. As such, we employ a lightweight slot-based token compression method to decrease the number of visual tokens to a fixed length, enabling video LLMs to sample more video frames. Numerical results illustrate the superior performance of the VTG-LLM compared to state-of-the-art video LLM methods. Additionally, ablation studies reveal that absolute-time tokens and sequence-time embedding significantly enhance the model’s performance in accurately locating timestamps. Furthermore, the slot-based token compression method outperforms the naive token compression baselines. Our key contributions are summarized as follows:
Figure 1: Overview of the VTG-LLM model. We utilize a fixed visual encoder and a Qformer model to encode sampled video frames into visual tokens. Subsequently, the absolute time $t_{i}$ and sequence ID $f_{i}$ are embedded and added to the top of each visual token for frame $i$. The slot-based token compression is designed to compress the number of tokens to a fixed number of visual tokens. To prevent overfitting and facilitate training, we employ a variety of prompts and well-structured answers. The timestamps with green color are encoded using six distinct time tokens.
  • We introduce VTG-IT-120K, a high-quality instruction tuning dataset comprising annotations for four distinct VTG tasks: moment retrieval (63.2K), dense video captioning (37.2K), video summarization (15.2K), and video highlight detection (3.9K).
  • We present a carefully crafted model, VTG-LLM, which incorporates timestamp information into video LLMs, providing models with the ability to understand timestamps. Quantitative results demonstrate the superior performance of VTG-LLM across various datasets, such as Charades-STA, QVHighlights, and YouCook2. Moreover, all three components are necessary for VTG-LLM to achieve the best results.
2 Related Works
@@
Video Temporal Grounding (VTG) aims to accurately locate the timestamps of events within a given video (Lin et al., 2023b). This encompasses tasks such as moment retrieval (Gao et al., 2017; Zala et al., 2023b; Oncescu et al., 2021; Hendricks et al., 2018a), dense video captioning (Zellers et al., 2021; Zala et al., 2023b; Tang et al., 2019; Fabian Caba Heilbron & Niebles, 2015), video summarization (Song et al., 2015; Gygli et al., 2014), and video highlight detection (Lei et al., 2021; Xiao et al., 2023). Traditional methods mainly address VTG tasks through large-scale video-text pre-training, using training objectives such as video-text contrastive learning (Xu et al., 2021; Wang et al., 2022; Yan et al., 2022), video-text matching (Li et al., 2023d; Chen et al., 2024), and masked auto-encoding (Tong et al., 2022; Zhao et al., 2024). Although these methods have shown satisfactory results, they require resource-consuming pre-training, lack zero-shot capabilities, and often need further fine-tuning on many downstream tasks.
@@
Recently, Video Large Language Models (video LLMs) have shown promise in video understanding tasks. Traditional video LLMs mainly generate captions that summarize videos (Zhang et al., 2023; Lin et al., 2023a; Li et al., 2023b; Maaz et al., 2023; Zhu et al., 2023). However, these models lack the capacity to accurately locate the timestamps of events within a given video. Some studies have attempted to address this issue. For example, TimeChat (Ren et al., 2023) constructs time-sensitive instruction tuning datasets and encodes timestamp knowledge into visual tokens. VTimeLLM (Huang et al., 2023) proposes a LLaVA-like three-stage training method. LITA (Huang et al., 2024) introduces fast-slow visual tokens and adds time tokens to LLM tokenizers. Momentor (Qian et al., 2024) proposes a time encoder to solve the quantization error of time tokens. HawkEye (Wang et al., 2024b) constructs a high-quality instruction tuning dataset based on InternVid (Wang et al., 2022) for the moment retrieval task. However, existing instruction tuning datasets for VTG tasks are either low-quality or exhibit extreme task imbalance, hindering model performance on overall tasks. Therefore, we propose a high-quality dataset VTG-IT to address this issue. Furthermore, we introduce VTG-LLM, which contains three well-designed components to boost video LLM performance on VTG tasks.
3 Method
@@
3.1 VTG-LLM: Incorporating Knowledge about Timestamps into Video LLMs
In this subsection, we introduce the detailed structure of VTG-LLM, which includes three key components designed to help video LLMs better understand timestamps. Specifically, we propose: (1) a sequence-time embedding mechanism to straight-forwardly incorporate timestamp information into visual tokens; (2) the introduction of absolute-time token without quantization error to distinguish time-related knowledge from other digit-related knowledge; and (3) the use of slot-based token compression to enable video LLMs to effectively handle more frames. We will introduce these components in detail in the following parts of this section.
3.1.1 Sequence-Time Embedding
@@
Existing studies (Zhang et al., 2023; Ren et al., 2023) employ sequence embedding to incorporate relative time information into visual tokens. Specially, given $N\times M$ tokens $\{\mathbf{z}_{i,j}|1\leq i\leq N,1\leq j\leq M\}$ sampled from $N$ frames, the process of adding sequence embedding onto visual tokens can then be represented by:
@@
Although promising, the sequence embedding only contains information about the temporal order and may not accurately represent knowledge about the absolute timestamps. For instance, sampled frames may not be uniformly distributed throughout the entire video (Ren et al., 2023; Huang et al., 2023), and the start time of the first sampled frame is also unknown. This makes it difficult to infer the timestamps of the frames. To address this issue, in addition to the sequence embedding, we also add the absolute time embedding as shown in the following equation
@@
Here, $t$ is the absolute timestamp (in seconds) of the corresponding frames. $\mathbf{W}_{t}\in\mathbb{R}^{T\times d}$ is the absolute time embedding. $T$ is the maximum timestamp (in seconds), which is set to 8192 by default, supporting videos for more than 1 hours long theoretically. It’s important to note that to avoid disrupting the visual tokens generated by pre-trained visual encoders and to speed up the convergence, $\mathbf{W}_{t}$ is initialized by setting all entries to zero. Without the zero-initialization method, the performance of VTG-LLM significantly degraded, as shown in the "TE Random Initialize" part of Table 2.
@@
Remark 3.1 (Discussion on existing techniques that integrating information of timestamps into visual tokens.).
In addition to incorporating time embeddings into visual tokens, there are other techniques for integrating timestamp information into visual tokens. For instance, adding text inputs into Qformer (Ren et al., 2023) and inserting time tokens before visual tokens (Hua et al., 2024). However, we believe that these techniques are orthogonal to the time embedding approach, and it is also possible to combine these approaches in future work.
3.1.2 Absolute-Time Tokens
The integration of special time tokens into the tokenizer has proven advantageous for video temporal grounding tasks, as demonstrated in several studies (Yang et al., 2023; Qian et al., 2024; Huang et al., 2024). However, the use of relative time tokens (frame ID) has revealed certain limitations. Firstly, there is the issue of quantization errors, which prevent the models from predicting timestamps outside the sampled frames. Secondly, the token embedding requires training from scratch, and any changes to the number of time tokens necessitate further training. As a result, our goal is to develop a new time token mechanism that can both represent fine-grained time without quantization errors and represent a broad range of timestamps without requiring additional effort when sampling varying numbers of frames.
@@ 3.1.2 Absolute-Time Tokens
To resolve the issue of quantization errors, we introduce the concept of absolute-time tokens. As depicted in Figure 1, we have incorporated eleven time tokens into the tokenizer. These consist of ten digit time tokens representing the digits from 0 to 9, and an additional token for a decimal point. All timestamps are represented using six time tokens. For instance, the time 120.5 seconds would be formatted as <TIME_ZERO><TIME_ONE><TIME_TWO><TIME_ZERO><TIME_DOT><TIME_FIVE>. This formulation allows the time tokens to handle videos more than 1 hours in length, with the precision remaining constant regardless of video length increases. Importantly, it is essential to format all timestamps using the same number of time tokens. As shown in the "Time Token not Formatted" section of Table 2, we found that maintaining a consistent format for all timestamps significantly improves model performance.
@@
Although using the absolute time token resolves the quantization error, we discovered that randomly initialized time tokens negatively impact the original token embedding space, preventing the LLMs from learning accurate knowledge about the time tokens. This resulting in poor performance of models, as shown in the "Token Embedding not Initialized" part of Table 2. To tackle this issue, our goal is to initialize the time-related knowledge using the digit-related knowledge, as the digit-related knowledge has already been well-trained during the LLM pretraining stage.
This is achieved by adjusting the initialization of the weights of token embedding and LLM prediction head. Specifically, we consider the token embedding matrix $\mathbf{W}_{token}$. When initializing the embedding for the time token "<TIME_ONE>", we align the embedding of this time token with the embedding of the token "1":
$[\mathbf{W}_{token}]_{\text{ID}(\text{'<TIME\_ONE>'}),j}=[\mathbf{W}_{token}]_{\text{ID}(\text{'1'}),j}\,,\forall 0\leq j<d\,,$ (5)
@@
3.1.3 Slot-Based Token Compression
Compressing the number of visual tokens is a critical issue in video LLMs. On one hand, using all the visual tokens generated by visual backbones preserves all the information. However, due to the restrictions on LLM context length, the number of sampled frames must also be reduced. On the other hand, sampling a subset of frames may result in overlooking important information. To resolve this issue, we propose a simple yet effective method, namely slot-based token compression, to compress the visual tokens to a fixed number of tokens.
(a) Overview of VTG-IT
(b) Slot-Based Token Compression
Figure 2: Overview of VTG-IT and slot based sampler. As shown in Figure 2(a), the VTG-IT contains annotations of 4 important video temporal grounding tasks. The VTG-IT-X indicate self-collected annotations. Figure 2(b) shows the architecture of slot-based token compression. The $N$ tokens are passed through a trainable slot dispatcher, which assigns dispatch weights $\bm{\omega}\in\mathbb{R}^{N}$ to each slot. The tokens are then mixed based on these dispatch weights to obtain the slots.
As illustrated in Figure 2(b), given $N$ visual tokens $\mathbf{z}_{1},\cdots,\mathbf{z}_{N}$, and the trainable slot dispatcher $\mathbf{\Phi}\in\mathbb{R}^{K\times d}$, where $K$ represents the number of slots, the visual tokens first pass through the slot dispatcher to obtain the dispatch weights, and the tokens are mixed based on the dispatch weights to generate the slots. For instance, the slot $k$ is generated by
@@ 3.1.3 Slot-Based Token Compression
Remark 3.2 (Discussion on slot mechanism).

### v3

@@ Abstract
Video Temporal Grounding (VTG) strives to accurately pinpoint event timestamps in a specific video using linguistic queries, significantly impacting downstream tasks like video browsing and editing. Unlike traditional task-specific models, Video Large Language Models (video LLMs) can handle multiple tasks concurrently in a zero-shot manner. Consequently, exploring the application of video LLMs for VTG tasks has become a burgeoning research area. However, despite considerable advancements in video content understanding, video LLMs often struggle to accurately pinpoint timestamps within videos, limiting their effectiveness in VTG tasks. To address this, we introduce VTG-LLM, a model designed to enhance video LLMs’ timestamp localization abilities. Our approach includes: (1) effectively integrating timestamp knowledge into visual tokens; (2) incorporating absolute-time tokens to manage timestamp knowledge without concept shifts; and (3) introducing a lightweight, high-performance, slot-based token compression technique designed to accommodate the demands of a large number of frames to be sampled for VTG tasks. Additionally, we present VTG-IT-120K, a collection of publicly available VTG datasets that we have re-annotated to improve upon low-quality annotations. Our comprehensive experiments demonstrate the superior performance of VTG-LLM in comparison to other video LLM methods across a variety of VTG tasks.
Codehttps://github.com/gyxxyg/VTG-LLM
Introduction
Figure 1: Overview of the VTG-LLM model.
@@ Introduction
Video Temporal Grounding (VTG) is a crucial component of video understanding. It requires models to accurately pinpoint event timestamps within a video based on the given query. This task is essential for subsequent operations such as video browsing and editing. In this paper, we adopt the categorization described by Lin et al. (2023b), and divide the VTG task into four primary sub-tasks. For instance, moment retrieval and dense video captioning tasks (Fabian Caba Heilbron and Niebles 2015; Zhou, Xu, and Corso 2018) require models to generate one or multiple timestamp intervals, each accompanied by captions. The video highlight detection task (Lei, Berg, and Bansal 2021) necessitates models to produce a prominent score curve, while the video summarization task (Song et al. 2015) demands that models output a series of timestamps, each associated with their corresponding video frames.
Despite significant efforts by traditional task-specific models designed for various VTG tasks (Yang et al. 2023; Luo et al. 2023a; Lei, Berg, and Bansal 2021), these methods are limited in their ability to (1) handle multiple VTG tasks simultaneously and (2) provide zero-shot capabilities on VTG tasks, which are crucial for real-world applications. As a remedy, recent research has begun exploring the use of video LLMs (Lin et al. 2023a; Zhang, Li, and Bing 2023; Li et al. 2023a) as generalists (Ren et al. 2023; Huang et al. 2023) for addressing VTG tasks due to their capacity to handle various tasks concurrently in a zero-shot manner. However, several persistent issues hinder the effectiveness of current video LLMs in understanding timestamp knowledge, which in turn affects their performance on VTG tasks:
  • Visual inputs should contain sufficient and accurate timestamp information to help models understand when the visual content occurs in the videos.
  • Concept shifts occur when varying input data produce identical output targets, potentially obscuring decision boundaries (Moreno-Torres et al. 2012). This issue arises when using shared token embedding and classification heads for all digit-related knowledge. For example, the number ’20’ can appear in both a counting context, such as ’There are 20 people,’ and a temporal context like ’From 20-30 seconds.’ Although these digits have different meanings in these situations, they are forced to share the same decision boundary, making classification more challenging.
  • VTG tasks necessitate sampling more frames compared to other video tasks, i.e., VQA tasks typically require only 8 or 16 frames (Lin et al. 2023a; Cheng et al. 2024). However, models cannot make reliable predictions on visual content that is not included in the sampled frames fed into language models. Therefore, due to the limited context length of LLMs, it is essential to develop effective token compression methods that enable the sampling of more frames.
To address these issues, in this paper, we present a novel video LLM model, VTG-LLM, comprising three components that efficiently integrate timestamp knowledge into video LLMs, thereby enhancing their performance on VTG tasks: (1) we introduce a sequence-time embedding method that integrates accurate timestamp data into visual tokens; (2) we incorporate absolute-time tokens that specifically handle timestamp knowledge without introducing quantization errors; (3) we employ a lightweight and high-performance slot-based token compression method to decrease the number of visual tokens to a fixed length, allowing video LLMs to sample more video frames.
In addition to VTG-LLM, we also discovered that existing datasets either suffer from low quality (Zellers et al. 2021) or exhibit significant task imbalance (Ren et al. 2023; Wang et al. 2024b). To address this issue, we introduce VTG-IT-120K, a dataset comprising 47.2K videos and 120K annotations. We collected this dataset from publicly available sources and re-annotated the low-quality captions using Gemini-1.5 Pro. VTG-IT-120K covers most mainstream VTG tasks, including moment retrieval (63.2K), dense video captioning (37.2K), video summarization (15.2K), and video highlight detection (3.9K). This dataset exhibits a more balanced task distribution compared to existing datasets (Ren et al. 2023; Wang et al. 2024b).
Numerical results illustrate the superior performance of the VTG-LLM compared to state-of-the-art video LLM methods. Additionally, ablation studies reveal that absolute-time tokens and sequence-time embedding significantly enhance the model’s performance in accurately locating timestamps. Furthermore, the slot-based token compression method outperforms the naive token compression baselines such as sampling and cross-attention in our studied cases. Our key contributions are summarized as follows:
  • We present VTG-LLM, a versatile model aims to capture all four VTG tasks. By effectively incorporating timestamp information into video LLMs, VTG-LLM equips these models with the capability to comprehend and process timestamps.Moreover, we have proposed carefully crafted initialization strategies for VTG-LLM to take advantage of pretrained video LLM weights.
  • We present VTG-IT-120K, a high-quality instruction tuning dataset consisting of 51.9K re-annotated data annotations with superior quality.
  • Quantitative results demonstrate the superior performance of VTG-LLM across various datasets, such as Charades-STA, QVHighlights, and YouCook2. Our code and datasets are provided in the Supplementary Material.
Related Works
@@
Video Temporal Grounding (VTG) aims to accurately locate the timestamps of events within a given video (Lin et al. 2023b). This encompasses tasks such as moment retrieval (Gao et al. 2017; Zala et al. 2023b; Oncescu et al. 2021; Wang et al. 2024c), dense video captioning (Zellers et al. 2021; Zala et al. 2023b; Tang et al. 2019; Fabian Caba Heilbron and Niebles 2015), video summarization (Song et al. 2015; Gygli et al. 2014), and video highlight detection (Lei, Berg, and Bansal 2021; Xiao et al. 2023). Traditional methods mainly address VTG tasks through large-scale video-text pre-training, using training objectives such as video-text contrastive learning (Xu et al. 2021; Wang et al. 2022), video-text matching (Li et al. 2023b; Chen et al. 2024), and masked auto-encoding (Tong et al. 2022; Zhao et al. 2024). Although these methods have shown satisfactory results, they require resource-consuming pre-training, lack zero-shot capabilities, and often need further fine-tuning on many downstream tasks.
@@
Large language models (LLMs) (Achiam et al. 2023; Touvron et al. 2023) have exhibited considerable potential in capturing knowledge and tackling real-world challenges using a zero-shot approach. Recently, research has explored integrating knowledge from other modalities, such as vision (Liu et al. 2024) and audio (Ghosal et al. 2023), to enhance the capabilities of LLMs. Within the visual modality, video large language models (video LLMs) have emerged as a significant research area (Lin et al. 2023a; Zhang, Li, and Bing 2023; Li et al. 2023a; Song et al. 2024a, b). Traditional video LLMs primarily generate captions that summarize videos (Zhang, Li, and Bing 2023; Lin et al. 2023a; Li et al. 2023a; Maaz et al. 2023; Zhu et al. 2023), but they struggle to accurately pinpoint event timestamps within videos. Some studies have attempted to address this limitation, such as TimeChat (Ren et al. 2023), which constructs time-sensitive instruction tuning datasets and encodes timestamp knowledge into visual tokens. VTimeLLM (Huang et al. 2023) proposes a LLaVA-like three-stage training method, while LITA (Huang et al. 2024) introduces fast-slow visual tokens and adds time tokens to LLM tokenizers. Momentor (Qian et al. 2024) proposes a time encoder to solve time token quantization errors, and HawkEye (Wang et al. 2024b) constructs a high-quality instruction tuning dataset based on InternVid (Wang et al. 2022) for the moment retrieval task. NumPro (Wu et al. 2024) add frame numbers to frames for ease of temporal understanding. However, existing instruction tuning datasets for VTG tasks are often low-quality or exhibit extreme task imbalance, which hinders model performance across tasks. To address this issue, we propose a high-quality dataset, VTG-IT, and introduce VTG-LLM, a model comprising three well-designed components to enhance video LLM performance on VTG tasks.
Method
@@
Overview of VTG-LLM
In this subsection, we present the detailed structure of VTG-LLM, which comprises three key components designed to enhance video LLMs’ understanding of timestamps. Specifically, we propose: (1) a sequence-time embedding mechanism that directly incorporates timestamp information into visual tokens; (2) the introduction of absolute-time tokens without quantization errors to differentiate time-related knowledge from other digit-related knowledge; and (3) the implementation of slot-based token compression to enable video LLMs to effectively process more frames. We will discuss these components in detail throughout this section.
Sequence-Time Embedding
@@
Existing studies (Zhang, Li, and Bing 2023; Ren et al. 2023) employ sequence embedding to incorporate relative time information into visual tokens. Specially, given $N\times M$ tokens $\{\mathbf{z}_{i,j}|1\leq i\leq N,1\leq j\leq M\}$ sampled from $N$ frames, the process of adding sequence embedding onto visual tokens can then be represented by:
@@
Although promising, the sequence embedding only contains information about the temporal order and may not accurately represent knowledge about the absolute timestamps. For instance, sampled frames may not be uniformly distributed throughout the entire video (Ren et al. 2023; Huang et al. 2023). Moreover, the sampling intervals can vary significantly for videos of different lengths. These issues make it difficult to infer the timestamps of the frames simply using sequence embedding. To address this issue, in addition to the sequence embedding, we also add the absolute time embedding as shown in the following equation
@@
Here, $t$ is the absolute timestamp (in seconds) of the corresponding frames. $\mathbf{W}_{t}\in\mathbb{R}^{T\times d}$ is the absolute time embedding. $T$ is the maximum timestamp (in seconds). It’s important to note that to avoid disrupting the visual tokens generated by pre-trained visual encoders and to speed up the convergence, $\mathbf{W}_{t}$ is initialized by setting all entries to zero. Since we employ pretrained vision encoder and Qformer modules, the performance of VTG-LLM is significantly degraded without the zero-initialization method. This is evident in the ”TE Random Initialize” section of Table 2.
@@
Please refer to Appendix (Guo et al. 2024) for ablation studies on using test-time interpolation.
Discussion on existing techniques that integrating information of timestamps into visual tokens.
In addition to incorporating time embeddings into visual tokens, there are other techniques for integrating timestamp information into visual tokens. For instance, adding text inputs into Qformer (Ren et al. 2023) and inserting time tokens before visual tokens (Hua et al. 2024). However, we believe that these techniques are orthogonal to the time embedding approach, and it is also possible to combine these approaches in future work.
Absolute-Time Tokens
The incorporation of unique time tokens in the tokenizer has shown benefits in video temporal grounding tasks, as evidenced by various studies (Yang et al. 2023; Qian et al. 2024; Huang et al. 2024). However, using relative time tokens (frame ID) exposes some limitations. First, quantization errors grow linearly with video length, complicating fine-grained predictions for longer videos. Second, training token embedding from scratch hinders leveraging pretrained video LLMs’ benefits. Our goal is to create a novel time token mechanism that accurately represents fine-grained timestamps and easily adapts to pretrained video LLMs.
@@
To resolve the issue of quantization errors, we introduce the concept of absolute-time tokens. As depicted in Figure 1, we have incorporated eleven time tokens into the tokenizer. These consist of ten digit time tokens representing the digits from 0 to 9, and an additional token for a decimal point. All timestamps are represented using six time tokens. For instance, the time 120.5 seconds would be formatted as $\langle t_{0}\rangle\langle t_{1}\rangle\langle t_{2}\rangle\langle t_{0}\rangle\langle t_{dot}\rangle\langle t_{5}\rangle$. This formulation allows the time tokens to handle videos more than 1 hours in length, with the precision remaining constant regardless of video length increases. Importantly, it is essential to format all timestamps using the same number of time tokens. As shown in the ”Time Token not Formatted” section of Table 2, we found that maintaining a consistent format for all timestamps significantly improves model performance.
@@
While using absolute time tokens eliminates quantization errors, we found that randomly initialized time tokens adversely affect the original token embedding space, hindering LLMs from learning precise time token knowledge. Consequently, this leads to subpar model performance, as demonstrated in the ”Token Embedding not Initialized” section of Table 2. To tackle this issue, our goal is to initialize the time-related knowledge using the digit-related knowledge, as the digit-related knowledge has already been well-trained during the LLM pretraining stage.
This is achieved by adjusting the initialization of the weights of token embedding and LLM prediction head. For example, consider the token embedding matrix $\mathbf{W}_{token}$, when initializing the embedding for the time token ”$\langle t_{1}\rangle$”, we align the embedding of this time token with the embedding of the token ”1”:
$[\mathbf{W}_{token}]_{\text{ID}(\text{'$\langle t_{1}\rangle$'}),j}=[\mathbf{W}_{token}]_{\text{ID}(\text{'1'}),j}\,,\forall 0\leq j<d\,,$ (5)
@@
Slot-Based Token Compression
Compressing visual tokens is crucial in VTG tasks. On one hand, models cannot produce reliable predictions without adequate visual input. On the other hand, the context length of LLMs inherently imposes a limitation. To address this, we propose a straightforward yet efficient approach called slot-based token compression, which compresses visual tokens to a fixed number, thereby enabling models to sample more frames.
Figure 2: Overview of slot based token compression.
As illustrated in Figure 2, given $N$ visual tokens $\mathbf{z}_{1},\cdots,\mathbf{z}_{N}$, and the trainable slot dispatcher $\mathbf{\Phi}\in\mathbb{R}^{K\times d}$, where $K$ represents the number of slots, the visual tokens first pass through the slot dispatcher to obtain the dispatch weights, and the tokens are mixed based on the dispatch weights to generate the slots. For instance, the slot $k$ is generated by
@@
Discussion on slot mechanism